24 June 2013 LBC IQ Technical Observing Night 3

OSA: Steve Allanson
Telescope Manager: Kevin Newton
Observers: Steven Villanueva, Olga Kuhn

Summary

High winds and clouds delayed opening. Clouds lingered a little while but then moved off, and a second set, but not so thick, came through later.

Started the night with a short series of extra- and intra-focal pupils at a few different filters: as well as V-BESSEL on LBC-B, SDT_Uspec and g-SLOAN. The fuzzy edge of the blue pupil is seen in all.

Next moved on to collect another series of data for collimation. Like last night, we started at high elevation, went down to 15 deg, and then back up. I changed the seeing threshold for Z11,Z22 from 2.5 to 1.7" before-hand. Still, at very low elevations, we saw negative z4 and positive z11, z22 corrections being sent by FPIA on both sides. On blue, because the estimated seeing exceeded the threshold, no spherical corrections were sent and convergence was reached after a few iterations, though Z4 built up as it waited for red convergence. On red, estimated seeing was not below the limit and much like last night, many iterations were needed, and Z4, Z11 and Z22 built up. Plots of the mirror X,Y,Z,RX,RY positions (attached) appear similar to those of last night.

Had to close for high winds. Reopened in time to obtain a few extra/intra focal pupil images of smaller size, D=47pix.

Log

High wind and clouds at sunset. Delaying opening for these.

Our plan tonight is to:
  • collect more data for the collimation model, going down in elevation and then up.
    • We'll stay about 30min W of meridian, but probably will stay in the N because of the wind.
    • Will plan not to clear active optics, but
    • Changed BadSeeingLimit in lbcfpia.cfg file from 2.5 to 1.7. At low elevation and poor seeing, we saw runaway z4,z11,z22.
  • measure filter focus offsets. We are choosing two relatively dense star fields for this, one at 45 deg with relatively unchanging elevation, and one at ~75 deg.

03:40 Closed for wind and clouds.

04:35 Opening.

04:46 Initial focus/collimation at 72 deg elevation.
  • Some cloud. dofpia iterated twice on blue, once on red. Z11 corrections not sent (1.9") and sent (1.7") but left some z11. Red did not.
 -1925      0      0      0      0   1100      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  6007      0      0      0      0         0      0    y  1.35  2.37                      /newdata/lbcr.20130624.044848.fits  r-SLOAN 
  1925      0      0      0      0  -1100      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -2100      0      0         0      0   1200      0      y  0.00  0.00                     dofpia_sph3_adjustment             
  9728      0      0         0      0      0         0      y  1.97  3.83                      /newdata/lbcb.20130624.044848.fits  V-BESSEL
   537   -302   -218    190    0   -126     -35    y  1.70  0.28                      /newdata/lbcb.20130624.045026.fits  V-BESSEL
  1215      0      0         0      0    -694      0      y  0.00  0.00                     dofpia_sph3_adjustment             
  • Now that clouds have moved off and we see more pupils, we asked Steve A to clear active optics.

04:43 Run dofpia again.
  • converged with 1.34" seeing on Blue and 0.78" seeing on Red. Skirt is clearly visible.
    • RB_VV0exin.ob to take 4-chip extra/intra focal pupils at VV centered on 0mm for both sides.
    • RB_SY0exin.ob to do the same at S and Y. (note S best-focus is ~0.4 so centering on 0 is not the best thing).
    • RB_VV0exin
05:30 Run dofpia again.
  • RB_VV0, RB_gi (no 0), RB_VV0, RB_SY (no 0), RB_VV0.
    • extra-focal g lbcb.20130624.052936.fits
    • extra-focal V lbcb.20130624.053237.fits
    • extra-focal S lbcb.20130624.053539.fits
    • extra-focal V lbcb.20130624.053848.fits
  • all blue pupils: S, V, g, show 'skirt'. Tonight it appears asymmetric. The sharpest edges are at the base of the swing arm to the MV area, then it looks fuzzy around the rest.

Will start to collect data for the collimation model, as last night. Starting at high elevation, going down, and then back up again.
  • Steve A starting pointing log.
elev 83
  • deltIE = 11" and deltCA = 15"
  • second copoint pair lbcb.20130624.055723.fits and lbcr.20130624.055852.fits
    • blue IQ is pretty good (fwhm gradient)
elev 78
  • Note: mirrors appear to be flat. Ambient temperature also has been pretty flat for last ~30min.
  • second copoint pair lbcb.20130624.060926.fits and lbcr.20130624.060920.fits
    • blue images appear a little elongated upper left to lower right.
elev 67
  • second copoint pair lbcb.20130624.062041.fits and lbcr.20130624.062037.fits
    • blue image fwhm is larger (seeing worse down here, about 1", 5.1 pix) and elongated.
elev 60
  • good convergence here. Blue est seeing 1.4, Red est seeing 0.8".
  • second copoint pair: lbcb.20130624.063014.fits and lbcr.20130624.063010.fits
elev 55
  • on first copoint pair, blue image did not appear in /newdata (match to r 063621). Repeating OB.
  • lbcrangebal crashed because iraf/ds9 pipe got messed up. We've seen this before, just noting it here. Had to restart iraf and ds9.
  • second copoint pair: lbcb.20130624.064510.fits and lbcr.20130624.064509.fits
elev 50
  • on first copoint pair, notice both blue and red images appear elongated. See b 065026 and r 065023
  • second copoint pair: lbcb.20130624.065244.fits and lbcr.20130624.065240.fits.
    • actually, I think the images on-axis for the second pair look rounder. How does lbcrangebal/coma-free pointing change elongation? If image is already astigmatic?
elev 45
  • second copoint pair: lbcb.20130624.070318.fits and lbcr.20130624.070312.fits
    • elongation on both blue and red.
elev 35
  • clouds so dofpia found no pupils on first try. On second try, found many and converged on first iteration (1.3/0.9 est seeing).
  • second copoint pair: lbcb.20130624.071414.fits and lbcr.20130624.071410.fits
elev 25
  • red converged in 2 iterations at est seeing 1.2", blue est seeing 1.5". Spherical was starting to build up on blue, but did not continue.
 -1050      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -1589    -96    -59   -244    -81    191     -3    y  1.41  0.65                      /newdata/lbcb.20130624.071807.fits  V-BESSEL
  -780   -136     17    -74    -56    139      8    y  1.47  0.33                      /newdata/lbcb.20130624.071943.fits  V-BESSEL
  -211    -90    -87    -32    -71     55    -24    y  1.53  0.13                      /newdata/lbcb.20130624.072113.fits  V-BESSEL
  1112      0      0      0      0   -635      0    y  0.00  0.00                     dofpia_sph3_adjustment             
-875      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -1074   -111     38    -14    254    162      5    y  1.12  0.46                      /newdata/lbcr.20130624.071803.fits  r-SLOAN 
  -231    -66    153     57     89     89      8    y  1.23  0.15                      /newdata/lbcr.20130624.071939.fits  r-SLOAN 
  -486     35     11    -27      9    102     -1    y  1.16  0.22                      /newdata/lbcr.20130624.072110.fits  r-SLOAN 
   840      0      0      0      0   -480      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  • second copoint pair lbcb.20130624.072611.fits and lbcr.20130624.072608.fits
elev 15
  • Red est seeing is less than the 1.7" threshold. dofpia is in a z4/z11/z22 runaway on the red side. But on the blue side, seeing is greater than the 1.7" threshold and dofpia converged in 3 iterations.
  • Both blue and red pupils have a fuzzy edge.
  • There is also a temperature gradient across both primaries.
  • DIMM says 1.5" seeing
 -1050      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -965      2   -214    155   -210      0      0    y  1.87  0.41                      /newdata/lbcb.20130624.072826.fits  V-BESSEL
    54    160   -130    227    -56      0      0    y  1.98  0.15                      /newdata/lbcb.20130624.072957.fits  V-BESSEL
  -309     76    -87   -132    -14      0      0    y  1.89  0.16                      /newdata/lbcb.20130624.073129.fits  V-BESSEL
  -109     22   -221    -72     64      0      0    y  1.96  0.12                      /newdata/lbcb.20130624.073304.fits  V-BESSEL
   -42     31   -133    131     13      0      0    y  1.86  0.11                      /newdata/lbcb.20130624.073435.fits  V-BESSEL
  -566    211     90    167     -8      0      0    y  2.09  0.26                      /newdata/lbcb.20130624.073606.fits  V-BESSEL
   243   -110     64   -201     26      0      0    y  2.05  0.16                      /newdata/lbcb.20130624.073736.fits  V-BESSEL
  1050      0      0      0      0   -600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -875      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -1511    -16    -66   -101    176    264     39    y  1.64  0.63                      /newdata/lbcr.20130624.072821.fits  r-SLOAN 
  -910    -86    200    366     47    146     32    y  1.55  0.42                      /newdata/lbcr.20130624.072953.fits  r-SLOAN 
  -907     -2     75     -1    -24    193     26    y  1.56  0.39                      /newdata/lbcr.20130624.073126.fits  r-SLOAN 
  -889    101     78   -131     71    171     37    y  1.69  0.38                      /newdata/lbcr.20130624.073300.fits  r-SLOAN 
  -983   -117    -47      5    -10    192     28    y  1.51  0.42                      /newdata/lbcr.20130624.073429.fits  r-SLOAN 
  -861    -82    255   -161    -67      0      0    y  1.80  0.37                      /newdata/lbcr.20130624.073603.fits  r-SLOAN 
   383     32   -109    -29     22      0      0    y  1.74  0.17                      /newdata/lbcr.20130624.073733.fits  r-SLOAN 
  1195      0      0      0      0   -683      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  • Note that the blue side, even though it did not send z11,z22 still send a lot of negative z4.
  • second copoint pair lbcb.20130624.074220.fits and lbcr.20130624.074214.fits
elev 25
  • Note now that after the first iteration on blue, a large positive z4 is sent to counter the negative. The red side also sent a large positive z4, but it did not send enough z11,z22 to remove the spherical put in.
 -1050      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  1923      9    209    -43    396      0      0    y  1.72  0.79                      /newdata/lbcb.20130624.074500.fits  V-BESSEL
  -339     -9    111   -100     -5    200     26    y  1.63  0.21                      /newdata/lbcb.20130624.074635.fits  V-BESSEL
  1170      0      0      0      0   -669      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -875      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  2622     36    -75     64   -284    -83    -12    y  1.32  1.05                      /newdata/lbcr.20130624.074454.fits  r-SLOAN 
   251    -25    -50     72     15    -25      7    y  1.35  0.13                      /newdata/lbcr.20130624.074629.fits  r-SLOAN 
   956      0      0      0      0   -546      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  • Red IQ is worse than Blue IQ on first copoint pair (see b 074902 and r 074858). Presumably because of so much spherical on red side.
  • second copoint pair is lbcb.20130624.075105.fits and lbcr.20130624.075102.fits
elev 35
 -1050     0      0      0        0       600      0     y  0.00  0.00                     dofpia_sph3_adjustment             
  1949    111   62    207    207   -132   -36    y  1.27  0.79                      /newdata/lbcb.20130624.075507.fits  V-BESSEL
   134    135   42     11      9        -9       13    y  1.42  0.10                      /newdata/lbcb.20130624.075640.fits  V-BESSEL
   -51     84      0      7      26        8       13     y  1.42  0.09                      /newdata/lbcb.20130624.075813.fits  V-BESSEL
  1042      0     0      0        0      -595      0     y  0.00  0.00                     dofpia_sph3_adjustment             
 -875      0      0      0      0       500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  1767  109   -52   302  -218  -283    -71  y  1.00  0.74                      /newdata/lbcr.20130624.075504.fits  r-SLOAN 
   662     -4   -125  -10   -197   -96     -5    y  1.20  0.30                      /newdata/lbcr.20130624.075637.fits  r-SLOAN 
   298     46    -69   89    135  -134    -14   y  1.12  0.18                      /newdata/lbcr.20130624.075809.fits  r-SLOAN 
   814      0      0      0      0     -465      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  • second copoint pair lbcb.20130624.080325.fits and lbcr.20130624.080321.fits
    • image quality better on blue than on red.
elev 45
  • second copoint pair lbcb.20130624.081325.fits and lbcr.20130624.081320.fits
elev 50
  • second copoint pair lbcb.20130624.082152.fits and lbcr.20130624.082148.fits
elev 55
  • catalog server went down --- we spent some time to use the USNO-B1.0 catalog web search form and create the OB.
    • we had been using a script which would launch ds9 and overplot the USNO-B1.0 catalog stars, but the catalog was not loading... it seems the connection to the catalog server is slow or flaky. Sometimes it works and sometimes not. We used this script all last night with no problem --- just started to have problems around 08:30 or 08:45 UT.
  • running dofpia.
  • second copoint pair lbcb.20130624.085239.fits and lbcr.20130624.085235.fits
elev 60
  • second copoint pair is lbcb.20130624.090656.fits and lbcr.20130624.090652.fits
    • elongation in both blue and red images.
elev 65
  • second copoint pair is lbcb.20130624.091346.fits and lbcr.20130624.091342.fits
elev 75
  • dofpia - FPIA complained that it could not compute sky, too many iterations (b 091733 r 091730)
  • second copoint pair is lbcb.20130624.092505.fits and lbcr.20130624.092502.fits
elev 85
  • slewed first to south, wind gusting from S and strong...
  • slewed back to north, but wind is just too strong, and we have to close.

09:35 UT closed due to high winds.

Attaching plots sxraw_20130624.jpg and dxraw_20130624.jpg to plot the mirror positions along the down/up tracks tonight. These look quite similar to the plots of last night. The temperature curves are for the same temperatures as last night: 9(blue), 10.75(black), 12.5(red). Tonight's spread in temperatures was less: T301 dropped from 11.8 to 10.4 deg C.

11:15 UT wind speeds dropped. Reopening to take some extra/intra focal pupils at a few elevations.

Steve A cleared active optics 77 deg
  • dofpia without /x2
  • RB_rV0ex4small to obtain 4-chip extra-focal pupils with M1 at dZ = -0.55mm. (/home/lbto/Calib_OBs/00FOCUS/).
    • D=47 pixels.
  • RB_rV0exin4small
  • RB_rV0exin.ob (usual 62pix pupils).
    • saturated.

Tuesday afternoon

At the end of the night, 20130624, we had re-opened, cleared active optics, and recollimated at high elevation, ~75 deg. I plotted the mirrorz position from the header of the last dofpia, corrected for 0.8mm and for the final Z4 exit correction scaled to mm:
  • for blue, lbcb.20130624.112742.fits Z=mirrorz from header. Z' = Z + 0.8 + (992 nm)*(0.0379 mm/ 1000 nm)
  • for red, lbcr.20130624.112738.fits Z=mirrorz from header, Z' = Z + 0.8 + (804 nm)*(0.0379 mm/ 1000 nm)
Without removing the temperature term, the two points lie much closer to the upward Z curve than to the initial downward one. When the temperature term is removed, the DX point lies closer to the upward curve, but the SX one lies on the downward curve, very close to the beginning-of-night starting point. Since Steve had cleared active optics, I would have expected both SX and DX points to be close to the downward curve.

-- OlgaKuhn - 24 Jun 2013
I Attachment Action Size Date Who Comment
dxraw_20130624.jpgjpg dxraw_20130624.jpg manage 159 K 24 Jun 2013 - 10:05 OlgaKuhn DX raw X,Y,Z,RX,RY mirror positions
sxraw_20130624.jpgjpg sxraw_20130624.jpg manage 153 K 24 Jun 2013 - 10:05 OlgaKuhn SX raw X,Y,Z,RX,RY mirror positions
Topic revision: r3 - 25 Jun 2013, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback